OpenStudy (theislander):

How do I put 2 points into point-slope form ex- (-5,-4,) and (-7,4)

OpenStudy (whpalmer4):

You want the equation of the line through those two points, written in point-slope form?

OpenStudy (theislander):

yes please

OpenStudy (texaschic101):

you have to find the slope first

OpenStudy (whpalmer4):

Okay. You have two points, (-5, -4) and (-7, 4). We'll call those points (x1, y1) and (x2, y2). Doesn't matter which point is which, so I'll choose (x1, y1) = (-5,-4). Now, to write the equation in point-slope form, we need two things: 1) a point 2) the slope We can pick either point for #1. For #2, we need to find the slope of the line going through those points. Here's the formula for the slope (m): rise (y2 - y1) m = ------- = -------- run (x2 - x1) Basically, you just look at the left-most point, figure out how many points you go to the right (the run) and the up/down (the rise)

OpenStudy (whpalmer4):

when you have the slope and a point, just plug the numbers into the formula: y - y1 = m ( x - x1 ) where (x1, y1) is a point on the line, and m is the slope.

OpenStudy (theislander):

alright so the slope is -4

OpenStudy (texaschic101):

slope = (4 - (-4) / (-7 - (-5) (4 + 4) / (-7 + 5) 8/-2 slope is -8/2 = -4 correct. now use point slope formula y - y1 = m(x - x1) slope(m) = -4 you can choose either of your sets of points (-5,-4) x1 = -5 and y1 = -4 now sub y - (-4) = -4(x - (-5) y + 4 = -4(x + 5) ===> point slope form

OpenStudy (texaschic101):

does that make sense to you ?

OpenStudy (theislander):

makes tons of sense. thank you

OpenStudy (texaschic101):

your very welcome :)